# On the generalized eigenvalue problem in subspace-based excited state methods for quantum computers

### v1

@@ II.3 Eigenvalue problem vs generalized eigenvalue problem
Figure 1: Eigenstates computed using (a) QSE in the case of the low condition number of overlap matrix (condition number is 7.1) and (b) q-sc-EOM. The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots on the x-axis. The geometry of the H4 molecule is shown in the figure.
@@ III Method
All computations in this study are carried out at a sto-3G basis using Jordan Wigner mapping to represent orbitals using qubits. All tests are carried out using an 8-qubit problem of H4 molecule with different bond lengths to make a consistent analysis throughout this study. The UCCSD ansatz within the VQE algorithm was used in all calculations for the estimation of an accurate ground state, which was followed by QSE or q-sc-EOM methods. The ground-state energy was computed using an infinite number of shots (exact estimation) to ensure all the effects studied resulted from the excited state method. The initial state was taken as the Hartree-Fock state. We then proceeded to the excited-state calculation using QSE and q-sc-EOM. Qiskit [80], Pennylane and PySCF were used in generating the codes for all calculations, which are provided in Refs. [81, 82]. The codes were verified using a second implementation using the Pennylane software [83] and an implementation in the ADAPT-VQE code [84].
@@ IV Results and Discussion
In the following subsections, we will provide a comparative analysis between QSE and q-sc-EOM in various situations of the condition number of the overlap matrix in QSE. It should be noted that although we are only using the cases of QSE and q-sc-EOM, the general conclusions are valid for any subspace-based excited state method that relies on solving a generalized eigenvalue equation (or eigenvalue equation) with elements measured on a quantum computer. Although there are no strict definitions, we will use the term “low condition number” for any matrix condition number low enough that it has almost no effect on the resulting eigenvalues. In the H4 cases studied, we have classified any condition number below 102 in this range. We define “medium condition number” as the range in which the condition number is not small enough so that its effects on the results can be ignored, but not large enough that techniques such as thresholding can be used. In the case of an 8 qubit H4 problem, we believe that this range is roughly between 102 to 104. A “large condition number” is defined in cases where the generalized eigenvalue is significantly ill-conditioned and cannot be solved without the use of the thresholding technique. In the H4 cases studied, we have classified any condition number above 104 into this class. These ranges may change with the size of the matrices involved, but the general conclusions in this study should hold for systems of any size.
@@ IV.1 Low condition number case
Figure 2: Eigenstates computed using QSE in the case of increasing condition number of overlap matrix of QSE (condition number indicated on subfigures). The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots on the x-axis. The geometry of the linear H4 molecule is shown in the subfigures.
@@ IV.2 Medium condition number case
Figure 3: Eigenstates computed using q-sc-EOM method for the H4 molecular geometry with condition number of overlap matrix in QSE of 592.5. The solid lines, the spheres and the bars have the same meaning as in previous plots.
Figure 4: Eigenstates computed using modified QSE method with artificially exact overlap matrix for the H4 molecular geometry with condition number of overlap matrix in QSE of 592.5. The solid lines, the spheres and the bars have the same meaning as in previous plots.
@@
Figure 5: Error in lowest 5 excited states vs the number of shots for H4 molecular geometry with condition number 592.5 (geometry shown in the figure).
@@ IV.3 High condition number case: Use of thresholding
(a) QSE with thresholding
@@
Figure 6: Eigenstates computed using (a) QSE method with thresholding (condition number of overlap matrix is 2.05$\times$10-12) and (b) q-sc-EOM method (without the need for thresholding) for the linear H4 molecular geometry. The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots on the x-axis. The geometry of the linear H4 molecule is shown in the figure.
The excited states energies of linear H4 molecules with all bond lengths of 5Å using QSE and q-sc-EOM are plotted in Fig. 6. The condition number of the overlap matrix in QSE in this case is 2.05 x 10 12. Due to the very high condition number, the QSE generalized eigenvalue equation could not be solved for a finite number of shots. The standard method to address this problem is thresholding (as discussed in Ref. [77]). We used a threshold of 5x10-03 for 103 and 104 shots, and 1x10-02 for 500 shots. These thresholds were selected as they were the smallest numbers that were able to solve the singularity problem in the selected cases. This thresholding process was able to provide solutions to the generalized eigenvalue equation, but the resulting solution has the following problems:
(i) Some of the excited states are missing in the solution achieved. These states are the green and brown colored lines around -0.3 to -0.6 Hartree energy and light blue close to -1.1 Hartree energy in the Fig. 6.
(ii) Most of the degenerate states from the solution are missing.
These could be due to the loss of information associated with the process of thresholding where some small eigenvalues and associated eigenvectors of the overlap matrix are deleted, which restricts the subspace of QSE on which the Hamiltonian is projected. We note that the q-sc-EOM method provides all the eigenstates consistent with the other examples shown in this paper. It may appear that QSE with thresholding has higher accuracy at a smaller shot count, but that effect may be because thresholding reduces the matrix size significantly and it may be an artifact of solving the generalized eigenvalue equation for much smaller matrices. The accuracy of q-sc-EOM is consistent in all cases studied.
@@ V Conclusions
In this work, we provide a quantitative study of the resilience of subspace-based quantum excited state methods to statistical sampling errors on a quantum computer. We choose QSE as a representative method for the class of excited-state methods that require solving the generalized eigenvalue equation on a classical computer with matrix elements evaluated on a quantum computer. We compared this with the q-sc-EOM method, which has an orthonormal set of excited wavefunctions, resulting in a working equation that requires solving an eigenvalue equation instead of the generalized eigenvalue equation. We divided the possible scenarios into three cases: the low condition number of the overlap matrix in QSE, the medium condition number, and the high condition number.
@@
We find that the performance of QSE and q-sc-EOM are similar when the condition number of the overlap matrix in QSE is small. At medium condition numbers, increasingly high errors in eigenstates arise in QSE as the condition number increases. This is because the inversion of an overlap matrix in the presence of noise is error-prone when the condition number is high enough. We note that although accuracy can be improved in the case of medium condition number with an increase in shots (if only shot noise is taken into consideration), it will require an order of magnitude more shots to reach a similar accuracy to q-sc-EOM in the examples studied. At very large condition numbers, the matrix becomes singular, and no solution can be found without the use of the thresholding technique [77]. Although the thresholding technique was able to solve the generalized eigenvalue equation, it resulted in missing excited states from the solution. This could be a potential problem in future chemical studies using these methods. We show that the problems we pointed to in this work at medium and large condition numbers can be eliminated with the use of the q-sc-EOM method.
@@
This study highlights the fact that QSE and other diagonalization-based excited-state methods that require the solution of a generalized eigenvalue equation are sensitive to errors on quantum computers, while subspace methods that require a solution to the eigenvalue equation, such as q-sc-EOM, are more resilient alternatives. Since the condition number of the overlap matrix in methods such as QSE can vary widely depending on molecular geometries, such methods could have significant variation in accuracy depending on the system studied. There is no general solution to this problem in QSE that we are aware of at medium condition number cases. Although thresholding can help solve the generalized eigenvalue equation in very high-condition number cases, the problem is not completely solved because the solution achieved through thresholding misses some of the excited states. We have used only statistical sampling errors in this study, but we believe that the qualitative conclusions will be valid for other sources of noise in the measurement of matrix elements. We also note that similar problems can be expected in other ground and excited state quantum algorithms that require solving a generalized eigenvalue equation in the presence of errors, and this study highlights potential challenges in such methods.
VI Acknowledgement
@@
AA would like to thank Nick Mayhall, Diksha Dhawan, Vibin Abraham and Ashutosh Kumar for useful discussions and their comments on the manuscript. AA, BG and SPS acknowledge startup funding from the University of North Dakota. AA and PFK want to acknowledge NSF, award number 2427046, for support. All authors acknowledge Amazon Braket and UND Computational Research Center for computing resources.

### v2

@@ II.4.3 Practical shot-noise guidance
Figure 1: Eigenstates computed using (a) QSE in the case of the low condition number of the overlap matrix (condition number is 7.1) and (b) q-sc-EOM. The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots per matrix element on the x-axis. The geometry of the H4 molecule is shown in the figure. This figure illustrates that when the overlap matrix is well-conditioned ($\kappa(\mathbf{S})=7.1$), QSE and q-sc-EOM show comparable shot-noise convergence and variance, i.e., overlap-conditioning-driven error amplification is not dominant in this regime.
II.5 Computational cost
The cost of subspace excited state methods can be discussed in circuit depth and the additional number of shots needed after the estimation of the ground state (summarized in table 1). In terms of the circuit depth, there are no additional gates needed in the circuit in QSE and qEOM beyond the ground state circuit and a negligible number of additional gates are needed (up to 7 CNOT gates) in q-sc-EOM. A major limitation of subspace excited state methods, including qEOM, QSE and brute force q-sc-EOM, is the shot count scaling for measuring all the relevant matrix elements in the Hamiltonian ($\mathbf{H}$) and overlap matrices ($\mathbf{S}$). The additional shot count scaling in QSE, qEOM and brute force q-sc-EOM are all $O(N^{12})$, where $N$ is the number of orbitals, due to the numerous matrix elements. A major advantage in q-sc-EOM in this regard is that the quantum version of the Davidson algorithm, developed by Kim and Krylov in Ref. [33], brings down the cost dramatically to $O(N^{8})$ without any approximation at the cost of additional circuit depth. Additional methods could further reduce the measurement cost in the Hamiltonian expectation value estimation step, thereby lowering the overall resource requirements.
Table 1: Computational cost in terms of additional circuit depth (in terms of scaling of CNOT gates) and additional shot count scaling of the three subspace excited state methods discussed. NN is the generic variable for the number of orbitals.
Method Shot count scaling
QSE $O(N^{12})$
qEOM $O(N^{12})$
q-sc-EOM $O(N^{12})$
q-sc-EOM with Davidson $O(N^{8})$
@@ III Method
All computations in this study are carried out in the STO-3G basis using Jordan–Wigner mapping to represent orbitals on qubits. All tests are carried out on an 8-qubit problem for the H4 molecule with different bond lengths to enable a consistent analysis throughout this study. The UCCSD ansatz within the VQE algorithm was used in all calculations for the estimation of an accurate ground state, which was followed by QSE or q-sc-EOM methods. The ground-state energy was computed using an infinite number of shots (exact estimation) to ensure all the effects studied resulted from the excited state method. The initial state was taken as the Hartree-Fock state. The parameters of UCCSD were optimized in the exact simulator using the L-BFGS optimizer to reach the noiseless UCCSD ground state. We then proceeded to the excited-state calculation using QSE and q-sc-EOM.
The q-sc-EOM and QSE excited-state calculations are carried out with shot noise (sampling errors) included, corresponding to a fixed number of shots for each study.
PennyLane [14], PySCF [75], and Qiskit [1] were used to generate the codes for all calculations, which are provided in Refs. [42, 41]. The codes were verified using an independent implementation in the PennyLane software and an implementation in the ADAPT-VQE code [3].
@@ IV Results and Discussion
In the following subsections, we will provide a comparative analysis between QSE and q-sc-EOM in various situations of the condition number of the overlap matrix in QSE. It should be noted that although we are only using the cases of QSE and q-sc-EOM, the general conclusions are valid for any subspace-based excited state method that relies on solving a generalized eigenvalue equation (or eigenvalue equation) with elements measured on a quantum computer. Although there are no strict definitions, we will use the term “low condition number” for any matrix condition number low enough that it has almost no effect on the resulting eigenvalues. In the H4 cases studied, we have classified any condition number below 102 in this range. We define “medium condition number” as the range in which the condition number is not small enough so that its effects on the results can be ignored, but not large enough that techniques such as thresholding can be used. In the case of an 8-qubit H4 problem, we believe that this range is roughly between 102 to 104. A “large condition number” is defined in cases where the generalized eigenvalue is significantly ill-conditioned and cannot be solved without the use of the thresholding technique. In the H4 cases studied, we have classified any condition number above 104 into this class. These numerical ranges are system- and noise-level dependent; a more general diagnostic is provided in Section II.4.3 in terms of $\lambda_{\min}(\mathbf{S})$ relative to the expected sampling error. These ranges may change with the size of the matrices involved, but the general conclusions in this study should hold for systems of any size.
@@ IV.1 Low condition number case
Figure 2: Eigenstates computed using QSE in the case of increasing condition number of overlap matrix of QSE (condition number indicated on subfigures). The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots per matrix element on the x-axis. The geometry of the linear H4 molecule is shown in the subfigures. This figure highlights the rapid amplification of sampling-noise sensitivity in QSE as $\kappa(\mathbf{S})$ increases (from 121.5 to 592.5), consistent with the $1/\lambda_{\min}(\mathbf{S})$ error-amplification factor in Eq. (24) and the shot-scaling trend in Eq. (28).
@@ IV.2 Medium condition number case
Figure 3: Eigenstates computed using q-sc-EOM method for the H4 molecular geometry with condition number of overlap matrix in QSE of 592.5. The solid lines, the spheres and the bars have the same meaning as in previous plots. This figure illustrates that q-sc-EOM remains stable for the same geometry where QSE exhibits a moderately ill-conditioned overlap matrix ($\kappa(\mathbf{S})=592.5$), consistent with the absence of overlap-matrix inversion when $\mathbf{S}=\mathbf{I}$.
Figure 4: Eigenstates computed using modified QSE method with artificially exact overlap matrix for the H4 molecular geometry with condition number of overlap matrix in QSE of 592.5. The solid lines, the spheres and the bars have the same meaning as in previous plots. This figure shows that using an (artificially) exact overlap matrix substantially reduces QSE errors, supporting the interpretation that noise in $\mathbf{S}$ and its subsequent inversion are key contributors to the observed instability.
@@
Figure 5: Error in lowest 5 excited states vs the number of shots for H4 molecular geometry with condition number 592.5 (geometry shown in the figure). This figure emphasizes the resource implication of overlap conditioning: for $\kappa(\mathbf{S})=592.5$, QSE requires approximately an order-of-magnitude more shots than q-sc-EOM to achieve comparable accuracy for the lowest excited states, consistent with Eq. (28).
(a) QSE with thresholding
(b) q-sc-EOM
[…]
@@ IV.3 High condition number case: Use of thresholding
The excited-state energies of linear H4 molecules with all bond lengths of 5 Å computed using QSE and q-sc-EOM are plotted in Fig. 6. The condition number of the overlap matrix in QSE in this case is $2.05\times 10^{12}$. Due to this very high condition number, the QSE generalized eigenvalue equation could not be solved for a finite number of shots. The standard method to address this problem is thresholding (as discussed in Ref. [21]). We used a threshold of $5\times 10^{-3}$ for $10^{3}$ and $10^{4}$ shots, and $1\times 10^{-2}$ for 500 shots. These thresholds were selected because they were the smallest values that resolved the singularity in the selected cases. This thresholding process was able to provide solutions to the generalized eigenvalue equation, but the resulting solution has the following problems:
(i) (see Fig. 6) Some excited states are missing from the computed spectrum. These correspond to the green- and brown-colored lines around 0.3-0.3 to 0.6-0.6 Hartree and the light-blue line near 1.1-1.1 Hartree.
(ii) Most of the degenerate states in the solution are missing.
This is due to the loss of information during thresholding: small eigenvalues (and their associated eigenvectors) of the overlap matrix are removed, which restricts the QSE subspace onto which the Hamiltonian is projected. Thresholding should be interpreted not only as a numerical stabilization procedure, but also as an explicit reduction of the effective subspace (model space), which can compromise spectral completeness. We note that the q-sc-EOM method provides the full set of eigenstates, consistent with the other examples shown in this paper. It may appear that QSE with thresholding has higher accuracy at a lower shot count, but this effect may arise because thresholding significantly reduces the matrix size and may be an artifact of solving the generalized eigenvalue equation for much smaller matrices. The accuracy of q-sc-EOM is consistent in all cases studied.
IV.4 Case with NH3 system
Building upon the observations in the H4 systems studied above, we extend our analysis to a more chemically relevant molecule, planar NH3 as a case study. Using an active space of (4e,4o) and the STO-6G basis set, we observe a similar trend in the excited state energies as seen in the case of H4. As the N-H bond of planar NH3 is stretched to mimic bond-breaking scenarios, the condition number of the overlap matrix in QSE increased to 124.4 and 427.6. This increase leads to an increase in numerical instability that appears as large deviations in the mean and large standard deviations in the measured energies using QSE, as shown in Fig. 7(a) and Fig. 8(a), respectively. It should be noted that in this case, the large amplification of errors due to condition numbers is seen even at relatively small condition numbers compared with the case of H4. This also indicates that the impact of the condition number on numerical instabilities is difficult to predict, as different molecular systems may have different instabilities for states at the same condition number. In contrast, q-sc-EOM remained consistently accurate for both geometries shown in Fig. 7(b) and Fig. 8(b).
(a) QSE
@@ IV.4 Case with NH3 system
Figure 7: Eigenstates of NH3 computed using (a) QSE in the case of the condition number of the overlap matrix (condition number is 124.4) and (b) q-sc-EOM. The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots per matrix element on the x-axis. The geometry of the NH3 molecule is shown in the figure. This figure shows that the conditioning-driven instability trends persist in a chemically relevant system: QSE exhibits increased deviations/variance at $\kappa(\mathbf{S})=124.4$, while q-sc-EOM remains comparatively stable.
(a) QSE
(b) q-sc-EOM
Figure 8: Eigenstates of NH3 computed using (a) QSE in the case of the condition number of the overlap matrix (condition number is 427.6) and (b) q-sc-EOM. The solid lines represent the exact solution (corresponding to an infinite number of shots) while the spheres and bars represent the mean and variance of the data in 10 simulations using the number of shots per matrix element on the x-axis. The geometry of the NH3 molecule is shown in the figure. This figure shows that further increase in overlap-matrix conditioning ($\kappa(\mathbf{S})=427.6$) leads to larger QSE instability, while q-sc-EOM remains comparatively stable, highlighting the geometry- and system-dependence of QSE reliability under sampling noise.
@@ V Conclusions
In this work, we provide both analytical and numerical analyses of the resilience of subspace-based quantum excited state methods to statistical sampling errors on a quantum computer. We choose QSE as a representative method for the class of excited-state methods that require solving the generalized eigenvalue equation on a classical computer with matrix elements evaluated on a quantum computer. We compared this with the q-sc-EOM method, which has an orthonormal set of excited wavefunctions, resulting in a working equation that requires solving an eigenvalue equation instead of the generalized eigenvalue equation. We divided the possible scenarios into three cases: the low condition number of the overlap matrix in QSE, the medium condition number, and the high condition number.
@@
We find that the performance of QSE and q-sc-EOM are similar when the condition number of the overlap matrix in QSE is small. At medium condition numbers, increasingly high errors in eigenstates arise in QSE as the condition number increases. This is because the inversion of an overlap matrix in the presence of noise is error-prone when the condition number is high enough. We note that although accuracy can be improved in the case of medium condition number with an increase in shots (if only shot noise is taken into consideration), it will require an order of magnitude more shots to reach a similar accuracy to q-sc-EOM in the examples studied. At very large condition numbers, the matrix becomes singular, and no solution can be found without the use of the thresholding technique [21]. Although the thresholding technique was able to solve the generalized eigenvalue equation, it resulted in missing excited states from the solution. This could be a potential problem in future chemical studies using these methods. We show that the dominant instability mechanism associated with solving a generalized eigenvalue problem with an ill-conditioned overlap matrix—in particular, the noise-sensitive inversion of $\mathbf{S}$—can be avoided in q-sc-EOM, whose working equation is a standard eigenvalue problem with an analytically identity overlap matrix ($\mathbf{S}=\mathbf{I}$). Our analytical study aligns with the numerical findings, and we find that the errors in eigenvalues are magnified in the generalized eigenvalue equation compared with the standard eigenvalue equation. The error magnification depends on the magnitude of the eigenvalue and inversely depends on the minimum eigenvalue of the overlap matrix.
@@
This study highlights the fact that QSE and other diagonalization-based excited-state methods that require the solution of a generalized eigenvalue equation are sensitive to sampling errors on quantum computers, while subspace methods that require a solution to the eigenvalue equation, such as q-sc-EOM, are more resilient alternatives with respect to the overlap-conditioning-driven error amplification mechanism discussed here. The conclusions are expected to qualitatively hold for other sources of noise as well, since the condition number of the overlap matrix will magnify the impact of other sources of errors in a similar way (in line with our analytical results). Sampling errors are used as a representative example of errors in this study, as we can expect that sampling errors will always exist in Hamiltonian expectation value estimation, even on a perfect quantum computer. The condition number of the overlap matrix in methods such as QSE can vary widely depending on molecular geometries; such methods could have significant variation in accuracy depending on the system studied. There is no general solution to this problem in QSE that we are aware of in the medium-condition-number regime. Although thresholding can help solve the generalized eigenvalue equation in very high-condition number cases, the problem is not completely solved because the solution achieved through thresholding misses some of the excited states. We have used only statistical sampling errors in this study, but we believe that the qualitative conclusions will be valid for other sources of noise in the measurement of matrix elements.
For balance, we note that q-sc-EOM and related linear-response/EOM approaches also have standard limitations. (i) The accuracy of excitation energies depends on the quality of the underlying ground-state wavefunction. (ii) Truncation of the excitation/operator manifold can limit spectral completeness and quantitative accuracy, particularly for states with strong multi-configurational character. (iii) Unlike ground-state VQE, excited-state solutions obtained from these linear-response/EOM constructions do not generally satisfy a strict variational bound. (iv) The measurement cost of estimating the Hamiltonian projection matrix elements can remain substantial for larger systems, however, it can be reduced substantially by quantum Davidson-type approaches [33] with additional circuit depth. Moreover, while $\mathbf{S}=\mathbf{I}$ removes the overlap-inversion amplification mechanism, statistical sampling errors and hardware noise in the Hamiltonian matrix elements can still shift eigenvalues according to standard eigenvalue perturbation behavior.
Spectroscopic applications in molecular sciences require the excited states within a specific energy range to be precisely known. Missing states or states with a high degree of uncertainty can lead to mislabeling or incorrect understanding of the electronic structure of the molecule; therefore, methods with a theoretically exact overlap matrix, such as q-sc-EOM, may be advantageous for such applications when spectral completeness is critical, provided that the ground-state quality and chosen excitation manifold are adequate for the target spectrum.
We also note that similar problems can be expected in other ground and excited state quantum algorithms that require solving a generalized eigenvalue equation in the presence of errors, and this study highlights potential challenges in such methods.
VI Acknowledgment
@@
AA would like to thank Nick Mayhall, Diksha Dhawan, Vibin Abraham and Ashutosh Kumar for their useful discussions and their comments on the manuscript. AA, BG and SPS acknowledge startup funding from the University of North Dakota. AA and PFK would like to acknowledge the NSF, award number 2427046 and 2429752, for support. All authors acknowledge Amazon Braket and UND Computational Research Center for computing resources.